Ball 1, with a mass of 150 g and traveling at 12 m/s , collides head on with ball 2, which has a mass of 320 g and is initially at rest.

Part A:

What is the final velocity of the ball 1 if the collision is perfectly elastic?

Express your answer in meters per second.

Part B:

What is the final velocity of the ball 2 if the collision is perfectly elastic?

Express your answer in meters per second.

Part C:

What is the final velocity of the ball 1 if the collision is perfectly inelastic?

Express your answer in meters per second.

Part D:

What is the final velocity of the ball 2 if the collision is perfectly inelastic?

Express your answer in meters per second.
Answer #1

a)

v1 = ( m1 - m2) u1 / ( m1 + m2)

v1 = (150 - 320)* 12 / ( 150 + 320)

v1 = - 4.32 m/s

negative sign indicate that the velocity os m1 after collision is opposite to its initial direction

=====

b)

v2 = 2 m1 u1 / ( m1 + m2)

v2 = 2* 150* 12 / ( 150 + 320)

v2 = 7.66 m/s

======

c)

in inelastic Collison both object has same velocity

m1 v1 = ( m1 + m2) v

150* 12 = ( 150 + 320) v

v = 3.83 m/s

======

d)

v2 = 3.83 m/s

======

Comment in case any doubt, will reply for sure . Goodluck